OpenStudy (anonymous):

dentify the equations of the asymptotes of the hyperbola (y - 7)2 - 16(x + 1)2 = 64.

OpenStudy (jdoe0001):

\(\bf \cfrac{(y-h)^2}{a^2}-\cfrac{(x-h)^2}{b^2}=1\\ \quad \\ \quad \\ (y - 7)^2 - 16(x + 1)^2 = 64\implies \cfrac{(y-7)^2}{64}-\cfrac{(x-(-1))^2}{64}=1\\ \quad \\ \implies \cfrac{(y-7)^2}{8^2}-\cfrac{(x-(-1))^2}{8^2}=1\\ \quad \\ \quad \\ \textit{asymptotes are at }\quad y=k\pm\cfrac{a}{b}(x-h)\)

OpenStudy (jdoe0001):

hmm I kinda ate the 16...

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

thanks for helping

OpenStudy (jdoe0001):

\(\bf \cfrac{(y-h)^2}{a^2}-\cfrac{(x-h)^2}{b^2}=1\\ \quad \\ \quad \\ (y - 7)^2 - 16(x + 1)^2 = 64\implies \cfrac{(y-7)^2}{64}-\cfrac{16(x-(-1))^2}{64}=1\\ \quad \\ \implies \cfrac{(y-7)^2}{8^2}-\cfrac{(x-(-1))^2}{2^2}=1\\ \quad \\ \quad \\ \textit{asymptotes are at }\quad y=k\pm\cfrac{a}{b}(x-h)\)

OpenStudy (jdoe0001):

arg... yet another typo =) one sec

OpenStudy (anonymous):

k

OpenStudy (jdoe0001):

\(\bf \cfrac{(y-\color{red}{k})^2}{\color{red}{a}^2}-\cfrac{(x-\color{red}{h})^2}{\color{red}{b}^2}=1\\ \quad \\ \quad \\ (y - 7)^2 - 16(x + 1)^2 = 64\implies \cfrac{(y-7)^2}{64}-\cfrac{16(x-(-1))^2}{64}=1\\ \quad \\ \implies \cfrac{(y-7)^2}{8^2}-\cfrac{(x-(-1))^2}{2^2}=1\\ \quad \\ \quad \\ \textit{asymptotes are at }\quad y=k\pm\cfrac{a}{b}(x-h)\)

OpenStudy (jdoe0001):

anyhow there

OpenStudy (anonymous):

so just plugin for the aysmptotes right

OpenStudy (jdoe0001):

pretty much, yes